Logiweb dictionary of HMpeano

Up Help

0 0 HMpeano

1 0 Mendelson 3 .2 (a)
2 0 Mendelson 3 .2 (b)
3 0 Mendelson 3 .2 (b)^{R}
4 0 Mendelson 3 .2 (c)
5 0 Mendelson 3 .2 (c)^{R}
6 0 Mendelson 3 .2 (c)_{h}^{R}
7 0 Mendelson 3 .2 (d)
8 0 Mendelson 3 .2 (d)^{R}
9 0 Mendelson 3 .2 (d)_{h}^{R}
10 0 Mendelson 3 .2 (f )
11 0 Mendelson 3 .2 (f )_{0}
12 0 Mendelson 3 .2 (f )_{n}
13 0 Mendelson 3 .2 (f )(ṫ)
14 0 Mendelson 3 .2 (g)
15 0 Mendelson 3 .2 (g)_{0}
16 0 Mendelson 3 .2 (g)_{n}
17 0 Mendelson 3 .2 (g)_{h}(ṙ, ṫ)
18 0 Mendelson 3 .2 (h)
19 0 Mendelson 3 .2 (h)_{0}
20 0 Mendelson 3 .2 (h)_{n}

21 0 S6′_{h}

22 0 Lemma 1
23 0 Lemma 2

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-07-04.UTC:21:57:57.981341 = MJD-53555.TAI:21:58:29.981341 = LGT-4627231109981341e-6